# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ Abstract
Phonons, the quantized vibrational modes of a crystal lattice, are ubiquitous quasiparticles in solid-state systems. They play a central role in a wide range of physical phenomena, from thermal transport—as primary carriers of heat in insulators—to their involvement in symmetry-breaking orders such as charge density waves and superconductivity. Traditionally considered as spinless bosons, phonons have recently emerged as a fertile ground for exploring topological physics, spurred by the rapid development of topological band theory initially formulated for fermionic systems. It is now understood that the phonon eigenstates, characterized by their eigenvalues and eigenvectors, can carry nontrivial topological invariants such as Berry curvature and Chern numbers, despite the absence of spin or charge. This new understanding opens up avenues to investigate the interplay between lattice dynamics, topology, and chirality in bosonic systems. In this article, we provide a comprehensive review of recent theoretical and experimental advances in the field of topological and chiral phonons. We begin by introducing the foundational concepts, including the classification of phononic band structures, symmetry-protected topological phases, and the definition of topological invariants in bosonic systems. Special attention is given to the concept of phonon angular momentum and its fundametal connection to Weyl phonons in inversion-symmetry-breaking systems. We then discuss key experimental realizations of topological and chiral phonons across a variety of material platforms. Finally, we outline outstanding challenges and promising directions for future research, such as the role of topology in phonon-mediated quasiparticle interactions and the manipulation of phonon angular momentum in quantum technologies.
[…]
@@
In recent years, inspired by conceptual advances in topology and chirality of electronic quantum states—such as topological insulators Hasan and Kane (2010); Qi and Zhang (2010a); Moore (2010), superconductors Qi and Zhang (2011); Ando and Fu (2015); Sato and Ando (2017), and semimetals Yan and Felser (2017); Armitage et al. (2018); Lv et al. (2021)—there has been a renewed interest in understanding phonons and their interactions with electronic excitations. Fundamentally, bulk band topology is characterized by topological invariants, such as the Berry phase and Chern number, which are defined on single-particle Bloch wave functions through Berry connection and Berry curvature Thouless et al. (1982a); Berry (1984). Importantly, the framework of topological band theory is not exclusive to fermionic quasiparticles—it is equally applicable to bosonic excitations, such as phonons and magnons, provided the system adheres to periodicity and satisfies the conditions of Bloch’s theorem.
@@
A key milestone was achieved in 2008, when Haldane and Raghu extended topological band theory to photonic crystals, proposing an electromagnetic analogue of the quantum Hall effect Haldane and Raghu (2008); Raghu and Haldane (2008), thereby establishing the first realization of noninteracting topological states in a bosonic system. The notion of topological phonons was subsequently introduced by Prodan in 2009, who demonstrated that nontrivial phononic Chern numbers can give rise to topologically protected edge modes—providing a novel mechanism to understand the dynamical instability of microtubules Prodan and Prodan (2009a). Later theoretical studies revealed that a nonzero phononic Chern number can arise in systems with broken time-reversal symmetry Sheng et al. (2006); Zhang et al. (2010) or broken spatial inversion symmetry Zhang et al. (2018a). These foundational insights paved the way for experimental realizations of topological phononic states across a wide range of platforms, including acoustic metamaterials Ma et al. (2019a); Xue et al. (2022), complex mechanical lattices Süsstrunk and Huber (2016), and crystalline solids Zhang et al. (2018a); Miao et al. (2018); Zhang et al. (2020b); Li et al. (2021a); Liu et al. (2022); Zhang et al. (2023b); Li et al. (2023a).
Although the development of topological phononics was initially inspired by advances in topological electronics, it should not be viewed as a straightforward or trivial extension. From a theoretical standpoint, phonons and electrons are governed by fundamentally different equations of motion. Phonons—manifesting as lattice vibrations—obey Newton’s second law, which involves a second-order time derivative, whereas electrons follow the Schrödinger equation, which is first-order in time. This fundamental difference makes the formulation of geometric phases for phonons less straightforward than for electrons Berry (1984). On the practical side, phonons are the primary carriers of heat in semiconductors and insulators, where electronic contributions to thermal transport are minimal or absent. The unique advantages of topological phonon modes—such as robust, defect-immune edge or surface transport channels—open promising avenues for next-generation phononic devices. In particular, they hold potential for information processing applications that require efficient thermal management and rapid heat dissipation.
In parallel with the development of topological phonons, chiral phonons, which are circularly polarized phonons with finite angular momentum Schaack (1976); Rebane (1983); Lin et al. (1985); McLellan (1988); Bermudez et al. (2008); Kagan and Maksimov (2008); Zhang and Niu (2015); Zhu et al. (2018); Komiyama and Murakami (2021), have also emerged as a rapidly expanding field of research, attracting significant interest in both microelectronics and quantum information science. Traditionally, phonons have been considered spinless quasiparticles carrying zero angular momentum. However, recent theoretical and experimental breakthroughs have overturned this long-standing assumption, demonstrating that phonons can indeed possess a finite angular momentum and pseudo-angular momentum under certain symmetry conditions. These modes have been shown to drive a variety of exotic physical phenomena, including circularly polarized Raman scattering Ishito et al. (2023b); Zhang et al. (2023b); Ishito et al. (2023a); Zhang et al. (2025a, b); Che et al. (2024), circular infrared absorption Zhu et al. (2018); Yang et al. (2024b), nonreciprocal thermal conductivity Qin et al. (2012); Kasahara et al. (2018); Grissonnanche et al. (2019); Zhang et al. (2019c); Li et al. (2020d); Grissonnanche et al. (2020); Chen et al. (2020, 2022); Li et al. (2023b); Ohe et al. (2024), and even spin Seebeck effects Fu et al. (2019); Fransson (2023); Kim et al. (2023); Li et al. (2024); Yao and Murakami (2025). Furthermore, the studies of phonon wavefunctions, which describe collective atomic displacements in real space, have revealed deep connections between topological and chiral phononsZhang et al. (2023b, 2025b).
@@
That said, the interpretation of certain experimentally observed phenomena attributed to topological and chiral phonons remain the subject of active debate. In some cases, a careful re-evaluation is necessary to ensure that the interpretations are fully consistent with the underlying physical mechanisms. In this article, we aim to give an overview of theoretical and experimental progresses on topological and chiral phonons in crystalline materials, whose energy spectrum spans from zero to tens of Terahertz. The structure of this review is organized below: Section I will briefly overview theoretical and experimental fundations of phonons in crystalline materials. Section II will introduce topological band theory and topological phonons in 3D. Section III will focus on topological phonons in in low-dimensions. Topological phonons in realistic materials will be discussed in Section IV. Section V introduce phonon (pseudo-)angular momentum and consequences of the (pseudo-)angular momentum conservation. We will also introduce the distinctions between two key concepts for chiral phonons in detail, $e.g.$, angular momentum and pseudo-angular momentum. The fundamental connection between topological phonons and chiral phonons will also be discussed. Parallel progress in experimental explorations on topological phonons and chiral phonns will be selected reviewed in Sec. VI. We will close the article with theoretical and experimental perspectives on topological and chiral phonons in Sec. VII.
@@ I Phonons: Theoretical Basis and Experimental Probes
In this section, we describe non-interacting phonons within the framework of the harmonic approximation. Our focus is on physical observables that related to experimental measurements. We will introduce material examples in both Secs.  IV and V to illustrate the experimental observation on topological phonons and chiral phonons.
@@ I.1 A Brief Overview of Lattice Dynamics
Phonons are ubiquitous quasiparticles describing the collective motions of the underlying lattice. The lattice Hamiltonian can be written as the sum of kinetic and potential energy. For simplicity, here we consider a system with only one atom in the unit cell Kittel and McEuen (2018):
${\mathcal{H}}=\frac{1}{2}\sum_{i,\alpha}M\dot{u}_{i}^{\alpha}\dot{u}_{i}^{\alpha}+\frac{1}{2}\sum_{i,\alpha}\sum_{j,\beta}\frac{\partial^{2}\Psi}{\partial u_{i}^{\alpha}\partial u_{j}^{\beta}}\bigg{|}_{\bm{0}}u_{i}^{\alpha}u_{j}^{\beta},$ (1)
where $u_{i}^{\alpha}$ represents atomic displacement away from its time-averaged equilibrium position; $i$ and $\alpha=x,y,z$ are labels of unit cell and vibrating degree of freedom, respectively. $M$ is the atomic mass. The Taylor expansion of the potential energy, $\Psi(\bm{u})$ follows
$\begin{split}\Psi(\bm{u})=\Psi(\bm{0})+\sum_{i,\alpha}\Psi^{\alpha}_{i}u^{\alpha}_{i}+\frac{1}{2}\sum_{i,\alpha}\sum_{j,\beta}\Psi^{\alpha\beta}_{ij}u^{\alpha}_{i}u^{\beta}_{j}+O(\bm{u}^{3}),\end{split}$ (2)
where $\Psi(\bm{0})$ is the potential energy at zero atomic displacements. Since we assume the lattice is in the equilibrium condition, the atomic force acting on $i$-th atom should be zero, i.e. $\Psi^{\alpha}_{i}=-\frac{\partial\Psi}{\partial u_{i}^{\alpha}}\big{|}_{\bm{0}}=0$. Under the harmonic approximation, only the second-order force constants $\Psi^{\alpha\beta}_{ij}=\frac{\partial^{2}\Psi}{\partial u^{\alpha}_{i}\partial u^{\beta}_{j}}\big{|}_{\bm{0}}$ are considered.
@@
where $I_{3\times 3}$ is the identity matrix, and $D\bm{(q)}$ is the dynamical matrix. The matrix elements of $D\bm{(q)}$ can be written as:
$D_{\alpha\beta}(\bm{q})\equiv\frac{1}{M}\sum_{i}\Psi^{\alpha\beta}_{ij}e^{-i\bm{q}\cdot(\bm{R}_{i}-\bm{R}_{j})},$ (4)
where $\bm{R}_{i}$ is the position vector of the unit cell containing the $i$-th atom. Since we considered the system with only one atom in a unit cell, Eq. (3) gives rise to three orthogonal eigenvectors, $\bm{\epsilon}_{\bm{q}\sigma}$, and eigenvales, $\omega_{\sigma}(\bm{q})$ ($\sigma=1,2,3$). All three modes are acoustic phonons. The atomic dynamics are superposition of the 3 normal modes:
$\bm{u}_{i}^{\alpha}=\frac{1}{\sqrt{NM}}\sum_{\bm{q},\sigma}\bm{\epsilon}_{\bm{q}\sigma}^{\alpha}X_{\bm{q}\sigma}e^{i\bm{q}\cdot\bm{R}_{i}},$ (5)
where $X_{\bm{q}\sigma}$ is the normal coordinates. Equation (1) can then be written in the normal coordinates as:
$\begin{split}\mathcal{H}&=\frac{1}{2}\sum_{\bm{q}\sigma}[\dot{X}_{\bm{q}\sigma}^{*}\dot{X}_{\bm{q}\sigma}+\omega_{\sigma}^{2}(\bm{q})X_{\bm{q}\sigma}^{*}X_{\bm{q}\sigma}]\\ &=\frac{1}{2}\sum_{\bm{q}\sigma}[P_{\bm{q}\sigma}^{*}P_{\bm{q}\sigma}+\omega_{\sigma}^{2}(\bm{q})X_{\bm{q}\sigma}^{*}X_{\bm{q}\sigma}],\end{split}$ (6)
@@
$a^{{\dagger}}_{\bm{q}\sigma}=\sqrt{\frac{\omega_{\sigma}(\bm{q})}{2\hbar}}\left(X_{\bm{-q}\sigma}+\frac{P_{\bm{q}\sigma}}{i\omega_{\sigma}(\bm{q})}\right),$ (7)
$a_{\bm{q}\sigma}=\sqrt{\frac{\omega_{\sigma}(\bm{q})}{2\hbar}}\left(X_{\bm{q}\sigma}-\frac{P_{\bm{-q}\sigma}}{i\omega_{\sigma}(\bm{q})}\right),$ (8)
and using the commutation relation [$P_{\bm{q}\sigma}$, $X_{\bm{q}^{\prime}\sigma^{\prime}}$]=$-i\hbar\delta_{\sigma\sigma^{\prime}}\delta_{\bm{q}\bm{q}^{\prime}}$, one derives:
@@
The above formalism can be generalized to $n$-atoms within a unit cell. In this case, there will be $3n$ eigenvectors and eigenvalues. Among the $3n$ phonon modes, 3 of them are acoustic phonons and the rest $33n-$ 3 modes are optical phonons. The three acoustic modes have zero frequency at $\Gamma$ ($\bm{q}=0$) because of the acoustic sum rule $\sum_{j}\Psi^{\alpha\beta}_{ij}=0$ which results from the translational invariance of the whole lattice potential energy.

### v2

# Advances in Phonons: From Band Topology to Phonon Chirality

@@ Abstract
Phonons are ubiquitous quasiparticles in solid state systems describing the quantized vibrations of a crystal lattice. Phonons play a central role in a wide range of physical phenomena, from transport to symmetry-breaking orders, such as charge density waves and superconductivity. Traditionally treated as spin-0 bosons that obey Bose-Einstein statistics, phonons have recently emerged as a fertile ground for exploring topological physics, spurred by the rapid development of topological band theory initially formulated for fermionic systems. It is now understood that the phonon eigenstates, characterized by their eigenvalues and eigenvectors, can carry nontrivial topological invariants, including the Berry phase and Chern number. This new understanding opens up avenues to investigate the interplay between lattice dynamics, topology, and chirality in bosonic systems. In this article, we review recent theoretical and experimental advances in the field of topological phonons and circularly polarized phonons. We introduce foundational concepts, including the classification of phononic band structures, symmetry-protected topological phases, and the definition of topological invariants in bosonic systems. We emphasize the concept of phonon angular momentum and its fundamental connection to Weyl phonons in $\mathcal{PT}$-breaking systems. Key experimental progresses on topological and circularly polarized phonons are discussed. We also outline outstanding challenges and promising directions for future research, such as the role of topology in phonon-mediated quasiparticle interactions and the manipulation of phonon angular momentum for potential applications in quantum technologies.
In condensed matter physics, phonons are ubiquitous lattice excitations that play a key role in thermal properties of crystalline materials and critical for emergent electronic orders such superconductivity and charge density waves. In contrast to electrons, which offer a rich palette of intrinsic degrees of freedom such as charge, spin, and orbital character for manipulation (Fig. 1), phonons are fundamentally different. As bosonic quasiparticles devoid of spin, charge, and orbital angular momentum, they have traditionally been viewed as possessing limited intrinsic tunability. This fundamental disparity has historically confined phonons to a more passive role in many physical processes and applications, where their influence is often leveraged indirectly through their interactions with other particles. It is precisely this limitation that motivates the search for novel strategies to actively control phononic properties.
@@
In recent years, inspired by conceptual advances in topology and chirality of electronic quantum states, such as topological insulators hasan2010colloquium ; qi2010quantum ; moore2010birth , superconductors qi2011topological ; ando2015topological ; sato2017topological , and semimetals yan2017topological ; Armitage2018Jan ; lv2021experimental , there has been a renewed interest in understanding phonons and their interactions with electronic excitations through the lens of topology and chirality. Fundamentally, the topology of a quantum excitation is characterized by topological invariants, such as the Berry phase and Chern number. These invariants are defined as integrals of the Berry connection and Berry curvature over a closed loop and manifold, respectively Thouless1982Aug ; Berry1984Mar . While initially developed for electrons, this framework is not exclusive to fermionic quasiparticles, but equally applicable to bosonic excitations, such as phonons ans magnons, provided the system adheres to periodicity and satisfies the conditions of Bloch’s theorem.
@@
Key milestones were achieved in 2004 and 2008, respectively, when Onoda, Haldane and Raghu extended topological band theory to photonic crystals, proposing an electromagnetic analogue of the quantum Hall effect onoda2004hall ; Haldane2008Jan ; raghu2008analogs , thereby establishing the foundation for subsequent studies of noninteracting topological states in bosonic systems. The notion of topological phonons was subsequently introduced by Prodan in 2009, who demonstrated that nontrivial phononic Chern numbers can give rise to topologically protected edge modes—providing a novel mechanism to understand the dynamical instability of microtubules Prodan2009Dec . Later theoretical studies revealed that a nonzero phononic Chern number can arise in systems with broken time-reversal symmetry sheng2006theory ; Zhang2010Nov or broken spatial inversion symmetry zhang2018double . These foundational insights paved the way for experimental realizations of topological phononic states across a wide range of platforms, including acoustic metamaterials Ma2019Apr ; Xue2022Dec , complex mechanical lattices Susstrunk2016Aug , and crystalline solids zhang2018double ; miao2018observation ; zhang2020twofold ; li2021observation ; liu2022ubiquitous ; zhang2023weyl ; li2023direct . Although the development of topological phononics was initially inspired by advances in topological electronics, it should not be viewed as a straightforward or trivial extension. From a theoretical standpoint, phonons and electrons are governed by fundamentally different equations of motion. Phonons—manifesting as lattice vibrations—obey Newton’s second law, which involves a second-order time derivative, whereas electrons follow the Schrödinger equation, which is first-order in time. This fundamental difference makes the formulation of geometric phases for phonons less straightforward than for electrons Berry1984Mar . We will discuss this matter in details in Section II. On the other hand, topological phonons share key features with topological electrons, such as robust and defect-immune edge or surface transport channels qin2024diverse ; liu2020topological ; ma2019topological ; wang2022topological ; Xue2022Dec ; yang2024topological . In semiconductors and insulators, since phonons are the main heat carriers, topological phonons may hold potential for information processing applications that require efficient thermal management and rapid heat dissipation.
Currently, topological phonons are reported to exhibit superior advantages in low-loss phononic wave guide Cha2018Dec ; Xi2025Jun , enhanced thermoelectric performance singh2018topological , and memory devices based on delay line technology Hafezi2011Nov ; Zhang2018Mar etc. The Bosonic nature of phonons has two consequences: First, the transport coefficients like phonon Hall conductivity is usually not quantized even when the phononic Chern number is nonzero so that it is usually not easy to capture the topological feature of phonons by transport measurement; Second, without a Fermi surface, all topological phonon modes are in principle detectable in the whole phonon band structure which makes the topological phononic material system much more versatile than the electronic conterpart, especially in the context of phononic metamaterials Xue2022Dec ; liu2020topological .
In parallel with the development of topological phonons, circularly polarized phonons with finite angular momentum (also called “chiral phonons” in more recent lateratures) schaack1976observation ; rebane1983faraday ; lin1985study ; mclellan1988angular ; bermudez2008chirality ; kagan2008anomalous ; zhang2015chiral ; zhu2018observation ; komiyama2021universal are evolving rapidly and attracting significant interests in both microelectronics and quantum information sciences. Angular momentum is a fundamental conserved quantity arising from spatial rotational symmetry, as established by Noether’s theorem. This intrinsic connection means that the conservation of angular momentum is a direct consequence of a system’s isotropy. In the continuum limit, phonons become isotropic and the angular momentum of phonons is described as spin-1 quantity, leading some studies to equate it with phonon spin jones1973asymmetric ; longuet1980spin ; bliokh2022field ; bliokh2022elastic ; yang2023hybrid ; shi2019observation ; bliokh2017optical ; bliokh2019transverse ; 1961PS ; 1962PS ; 1962PSLevine . However, in periodic structures such as crystalline solids and metamaterials, the angular momentum of phonons is no longer quantized but can take arbitrary magnitudes. Thus, some concepts of phonons are similar to spin-1 bosons, but they are not in a rigorous sense. These varying interpretations and definitions of phonon spin are highly context-dependent. Therefore, when discussing the concept of phonon spin, it is essential to clarify the underlying assumptions and conditions. (A similar multifaceted concept is “chirality in phonons”, which will be addressed in Chapter V.)
For a along time, the angular momentum of phonons have been overlooked. Recent theoretical and experimental breakthroughs have shown that phonons can indeed possess a finite angular momentum and pseudo-angular momentum under certain symmetry conditions zhang2025thechirality . These modes are believed to drive a variety of exotic physical phenomena, including circularly polarized Raman scattering ishito2023truly ; zhang2023weyl ; ishito2023chiral ; zhang2025thechirality ; zhang2025weyl ; che2025magnetic , circular infrared absorption zhu2018observation ; yang2025inherent and emission liu2019valley ; li2019momentum ; he2020valley ; kong2022comprehensive , nonreciprocal thermal conductivity qin2012berry ; kasahara2018majorana ; grissonnanche2019giant ; zhang2019thermal ; li2020phonon ; grissonnanche2020chiral ; chen2020enhanced ; chen2022large ; li2023phonon ; ohe2024chirality , and even spin Seebeck effects fu2019spin ; fransson2023chiral ; kim2023chiral ; li2024chiral ; yao2025theory . Interestingly, it is realized that topological and circularly polarized phonons are deeply connected in the context of Weyl phonons zhang2023weyl ; zhang2025weyl . That said, the interpretation of certain experimentally observed phenomena attributed to topological and circularly polarized phonons remain the subject of active debate. In some cases, a careful re-evaluation is necessary to ensure that the interpretations are fully consistent with the underlying physical mechanisms.
@@
In this review, we aim to give an overview of theoretical and experimental progresses on topological and circularly polarized phonons in crystalline materials, whose energy spectrum spans from zero to tens of Terahertz. The structure of this review is organized below: Section I will briefly overview theoretical and experimental foundations of phonons in crystalline materials. Section II will introduce topological band theory and topological phonons in 3D. Section III will focus on topological phonons in low-dimensions. Topological phonons in realistic materials will be discussed in Section IV. Section V introduce phonon (pseudo-)angular momentum and consequences of the (pseudo-)angular momentum conservation. We will also introduce the distinctions between two key concepts for circularly polarized phonons in detail, $e.g.$, angular momentum and pseudo-angular momentum. The fundamental connection between topological phonons and circularly polarized phonons will also be discussed. Parallel progress in experimental explorations on topological phonons and circularly polarized phonons will be selectivelly reviewed in Sec. VI. We will close the article with theoretical and experimental perspectives on topological and circularly polarized phonons in Sec. VII.
@@ I Phonons: Theoretical Basis and Experimental Probes
In this section, we describe non-interacting phonons within the framework of harmonic approximation. We focus on physical observable that is related to experimental measurements.
@@ I.1 A Brief Overview of Lattice Dynamics
Phonons are ubiquitous quasiparticles describing the collective motions of the underlying lattice. The lattice Hamiltonian can be written as the sum of kinetic and potential energy. For simplicity, we start with a system with one atom in the unit cell kittel2018introduction :
${\mathcal{H}}=\frac{1}{2}\sum_{i,\alpha}M\dot{u}_{i}^{\alpha}\dot{u}_{i}^{\alpha}-\frac{1}{2}\sum_{i,\alpha}\sum_{j,\beta}\frac{\partial^{2}\Psi}{\partial u_{i}^{\alpha}\partial u_{j}^{\beta}}\bigg|_{\bm{0}}u_{i}^{\alpha}u_{j}^{\beta},$ (1)
where $u_{i}^{\alpha}$ represents atomic displacement away from its time-averaged equilibrium position. Index $i$ and $j$ mark unit cell, and $\alpha,\beta=x,y,z$ represent vibrating directions. $M$ is the atomic mass. The Taylor expansion of the potential energy, $\Psi(\bm{u})$ follows
$\begin{split}\Psi(\bm{u})=\Psi(\bm{0})+\sum_{i,\alpha}\Psi^{\alpha}_{\ell i}u^{\alpha}_{i}+\frac{1}{2}\sum_{i,\alpha}\sum_{j,\beta}\Psi^{\alpha\beta}_{ij}u^{\alpha}_{i}u^{\beta}_{j}+O(u^{3}),\end{split}$ (2)
where $\Psi(\bm{0})$ is the potential energy without atomic displacements. Since we assume the lattice is in equilibrium condition, the atomic force acting on $i$-th atom should be zero, i.e., $\Psi^{\alpha}_{i}=\frac{\partial\Psi}{\partial u_{i}^{\alpha}}\big|_{\bm{0}}=0$. Under the harmonic approximation, only the second-order force constants $\Psi^{\alpha\beta}_{ij}=\frac{\partial^{2}\Psi}{\partial u^{\alpha}_{i}\partial u^{\beta}_{j}}\big|_{\bm{0}}$ are considered. In the presence of transnational symmetry, $\Psi^{\alpha\beta}_{ij}=\Psi^{\alpha\beta}(\mathbf{R}_{i}-\mathbf{R}_{j})=\Psi^{\alpha\beta}_{l}(\mathbf{R}_{l})$. Here $\bm{R}_{l}=\bm{R}_{i}-\bm{R}_{j}$ is the relative distance between the $i$-th and $j$-th unit cell.
@@
where $I_{3\times 3}$ is the rank-3 identity matrix. The $D\bm{(q)}$ is the dynamical matrix, which element can be written as:
$D_{\alpha\beta}(\bm{q})\equiv\frac{1}{M}\sum_{l}\Psi^{\alpha\beta}_{l}e^{-i\bm{q}\cdot\bm{R}_{l}},$ (4)
Eq. (3) gives rise to three orthogonal eigenvectors, $\bm{\epsilon}_{\bm{q}\sigma}$, and eigenvales, $\omega_{\sigma}(\bm{q})$ ($\sigma=1,2,3$). All three modes are acoustic phonons.
For topological and circularly polarized phonons, one need to go beyond the single-atom framework. For a lattice with $r$-atoms in one unit cell, the dynamical matrix is generalized to:
$D_{\alpha\beta}\left(\begin{array}[]{c}\mathbf{q}\\ \mathbf{s,s^{\prime}}\end{array}\right)\equiv\frac{1}{\sqrt{M_{s}M_{s^{\prime}}}}\sum_{l}\Psi^{\alpha\beta}\left(\begin{array}[]{c}l\\ \mathbf{s,s^{\prime}}\end{array}\right)e^{-i\bm{q}\cdot\bm{R}_{l}},$ (5)
where $\Psi^{\alpha\beta}\left(\begin{array}[]{c}l\\ \mathbf{s,s^{\prime}}\end{array}\right)=\Psi^{\alpha\beta}\left(\begin{array}[]{c}i-j\\ \mathbf{s,s^{\prime}}\end{array}\right)=\frac{\partial^{2}\Psi}{\partial u^{\alpha}_{i}(s)\partial u^{\beta}_{j}(s^{\prime})}\big|_{\bm{0}}$. We use $\bm{s}$ and $\bm{s}^{\prime}$ to indes atoms in a unit cell. The eigen equation can then be generalized to:
$\mathrm{det}|D_{\alpha\beta}\left(\begin{array}[]{c}\mathbf{q}\\ \mathbf{s,s^{\prime}}\end{array}\right)-\omega^{2}\mathit{I}_{3r\times 3r}|=0.$ (6)
Eq. (6) gives rise to $3r$ orthogonal eigenvectors, $\bm{\epsilon}_{\bm{q}\sigma}$, and eigenvales, $\omega_{\sigma}(\bm{q})$ ($\sigma=1,...,3r$), corresponding to three acoustic modes and (3r33r-3) optical modes. The atomic distortion can be written as:
$\bm{u}_{i}^{\alpha}(s)=\frac{1}{\sqrt{NM_{s}}}\sum_{\bm{q},\sigma}\bm{\epsilon}_{\bm{q}\sigma}^{\alpha}(s)Q_{\bm{q}\sigma}e^{i\bm{q}\cdot\bm{R}_{i}},$ (7)
where $Q_{\bm{q}\sigma}$ is the normal coordinates. The multi-atom version of Eq. (1) can then be written in the normal coordinates as:
$\begin{split}\mathcal{H}&=\frac{1}{2}\sum_{\bm{q}\sigma}[\dot{Q}_{\bm{q}\sigma}^{*}\dot{Q}_{\bm{q}\sigma}+\omega_{\sigma}^{2}(\bm{q})Q_{\bm{q}\sigma}^{*}Q_{\bm{q}\sigma}]\\ &=\frac{1}{2}\sum_{\bm{q}\sigma}[P_{\bm{q}\sigma}^{*}P_{\bm{q}\sigma}+\omega_{\sigma}^{2}(\bm{q})Q_{\bm{q}\sigma}^{*}Q_{\bm{q}\sigma}],\end{split}$ (8)
@@
$a^{{\dagger}}_{\bm{q}\sigma}=\sqrt{\frac{\omega_{\sigma}(\bm{q})}{2\hbar}}\left(Q_{\bm{-q}\sigma}+\frac{P_{\bm{q}\sigma}}{i\omega_{\sigma}(\bm{q})}\right),$ (9)
$a_{\bm{q}\sigma}=\sqrt{\frac{\omega_{\sigma}(\bm{q})}{2\hbar}}\left(Q_{\bm{q}\sigma}-\frac{P_{\bm{-q}\sigma}}{i\omega_{\sigma}(\bm{q})}\right),$ (10)
and using the commutation relation $P_{\bm{q}\sigma}$, $Q_{\bm{q}^{\prime}\sigma^{\prime}}$$i\hbar\delta_{\sigma\sigma^{\prime}}\delta_{\bm{q}\bm{q}^{\prime}}$, one derives: